In the laboratory you dissolve 16.3 g of iron(III) bromide in a volumetric flask and add water to a total volume of 375 . mL.

What is the molarity of the solution?
1
Expert's answer
2020-09-29T06:55:41-0400

Iron II Bromide"= FeBr_3" Molar mass of= "295.56 g\/mol"


"moles = 16.3 g \times \frac{1 mol}{295.56 g} = 0.056moles"


liters of the solution = 375 ml x 1 L/1000 ml = 0.375 L


"Molarity=\frac{Moles}{Volume}=\frac{0.056}{0.375}=0.147M"